# Whispering in the dark

### v1

@@ Abstract
Context. Recently, astrometric and spectroscopic surveys of OB stars have revealed a few stellar-mass black holes (BHs) with orbital periods as low as 10 days. Contrary to wind-fed BH high-mass X-ray binaries, no X-ray counterpart has been detected, probably because of the absence of a radiatively efficient accretion disk around the BH. Yet, dissipative processes in the hot, dilute and strongly magnetized plasma around the BH (so-called BH corona) can still lead to non-thermal X-ray emission (e.g. synchrotron).
@@
Methods. We use detailed binary evolution models computed with MESA for initial primary masses of 10-90 and orbital periods from 1-3000 d. The X-ray luminosity is computed for a broad range of radiative efficiencies that depend on the mass accretion rate and flow geometry.
Results. For typical conditions around stellar-mass BHs, we show that particle acceleration through magnetic reconnection can heat the BH corona. A substantial fraction of the gravitational potential energy from the accreted plasma is converted into non-thermal X-ray emission. Our population synthesis analysis predicts at least 28 (up to 72) BH+OB star binaries in the Large Magellanic Cloud (LMC) to produce X-ray luminosity above 1031 erg s-1, observable through focused Chandra observations. We identify a population of SB1 systems in the LMC and HD96670 in the Milky Way comprising O stars with unseen companions of masses above 2.3 that aligns well with our predictions and may be interesting sources for follow-up observations. The predicted luminosities of the OB companions to these X-ray-emitting BHs are 104.5-5.5 .
Conclusions. These results make the case for long-time exposure in X-rays of the stellar-mass BH candidates identified around OB stars. It will constrain the underlying population of X-ray-faint BHs, the evolution from single to double degenerate binaries, and the progenitors of gravitational wave mergers.
@@
The detection of merging stellar-mass black holes (BHs) and neutron stars (Abbott et al. 2016, 2019) has ushered in a thrilling quest to discover their progenitors. Co-evolution of massive binary stars (i.e. hydrogen burning starts at the same instant of time in both binary components) is a possible formation channel of merging compact objects (Belczynski et al. 2008; Marchant et al. 2016; Kruckow et al. 2018; Spera et al. 2019; Vigna-Gómez et al. 2019; Mapelli 2020; Bavera et al. 2021; Marchant et al. 2021; Jiang et al. 2023). The stable mass transfer channel has been shown to significantly contribute to double compact object mergers (van den Heuvel et al. 2017; Gallegos-Garcia et al. 2021; van Son et al. 2022; Shao & Li 2022; Briel et al. 2023; Picco et al. 2024; Dorozsmai & Toonen 2024; Olejak et al. 2024). However, many unconstrained assumptions (such as mass and angular momentum loss, internal mixing, stellar winds) used in the modelling of binary evolution cripple our capacity to make reliable predictions regarding the rates and properties of double compact object mergers (Broekgaarden et al. 2022; Mandel & Broekgaarden 2022; Marchant & Bodensteiner 2023).
@@
In the co-evolutionary scenario, high-mass X-ray binaries (HMXBs) represent an ephemeral albeit decisive phase before the formation of double compact objects from massive binaries. HMXBs contain a massive star in orbit with an X-ray bright compact object (Walter et al. 2015; Kretschmar et al. 2019; Motta et al. 2021; Chaty 2022; Fortin et al. 2023). A few of them were found to host stellar-mass BHs accreting stellar material from mass transfer via Roche lobe overflow and/or the capture of a fraction of the stellar wind (e.g. Cygnus X-1 Orosz et al. 2011, and LMC X-1 Orosz et al. 2009). Beyond their role in explaining mergers, these systems are excellent testbeds for theories of stellar evolution (Belczynski et al. 2008; Fishbach & Kalogera 2022), gravity and magnetism in the strong field regime (Narayan & Yi 1995; Bozzo et al. 2008; Karino et al. 2019), stellar winds and X-ray irradiation (Krtička et al. 2018; Vilhu et al. 2021).
@@
The plasma surrounding accreting stellar-mass BHs is a source of X-rays whose intensity and spectral properties depend on the flow geometry (disk or sphere-like) and on the mass accretion rate (Shakura 1973; Narayan & Yi 1995; Frank et al. 2002). At high mass accretion rates, a geometrically thin and optically thick accretion disk can form which is radiatively efficient and X-ray bright. But at lower mass accretion rates (Sharma et al. 2007; Xie & Yuan 2012) and/or when the flow is not centrifugally supported (Narayan & Yi 1994), an accretion disk seldom forms and the radiative efficiency of the plasma drops. In very sub-Eddington accretion flows, most of the kinetic energy of the electrons is not radiated away but advected into the BH event horizon (ADAF, Ichimaru 1977; Narayan & Yi 1994; Narayan et al. 1998; Yuan 2001). Also, a significant fraction of the material captured by the gravitational potential of the BH might eventually not accrete onto the BH but is lost in strong winds (ADIOS – Blandford & Begelman 1999; Yuan et al. 2012, 2015). As a result, the X-ray luminosity of most stellar-mass BHs might be below the detectability threshold of all-sky instruments ( erg s-1, taken from Vanbeveren et al. 2020), although they can be detected during outbursts.
@@
Population synthesis models predict thousands of BHs in orbit with a main sequence, giant or supergiant companions in the Milky Way (Shao & Li 2019, 2020). Yet, most BHs in a binary with a main sequence OB star are expected to be X-ray faint due to the inability to form a radiatively efficient accretion disk around the BH (Sen et al. 2021; Hirai & Mandel 2021). In the absence of an accretion disk, X-rays are dominated by non-thermal emission from electrons accelerated in the hot, dilute, magnetized and spherical (i.e. non-centrifugally maintained) region around the BH, called the BH “corona” (Bisnovatyi-Kogan & Lovelace 1997; Quataert & Narayan 1999). This radiative model has been corroborated by observations of supermassive BHs like Sagittarius A∗ (Baganoff et al. 2003), M87∗ (Event Horizon Telescope Collaboration et al. 2019) and NGC 3115 (Wong et al. 2011). In this work, we use the above model to investigate the properties of X-ray emission from stellar-mass BHs in orbit with a main sequence OB star.
@@
The sample of detected stellar-mass BHs is likely very incomplete. The Milky Way and Magellanic Clouds are predicted to harbour millions and hundreds of stellar-mass BHs respectively (Timmes et al. 1996; Breivik et al. 2017; Langer et al. 2020; Chawla et al. 2022; Jayasinghe et al. 2023). Yet, only a few tens were confirmed through X-ray detection (Walter et al. 2015; Corral-Santana et al. 2016) and, more recently, a handful through monitoring of the stellar companion (Giesers et al. 2018; Thompson et al. 2019; Liu et al. 2019; Rivinius et al. 2020; Masuda & Hirano 2021; Jayasinghe et al. 2021; Gomez & Grindlay 2021; Saracino et al. 2022; Shenar et al. 2022a; Tanikawa et al. 2023; El-Badry et al. 2023a, b; Gaia Collaboration et al. 2024). Most of the 60 stellar-mass BHs detected through X-ray emission are transients, with periodicities of the order of years to decades (McClintock & Remillard 2006; Corral-Santana et al. 2016). In contrast, their X-ray luminosity during quiescence is several orders of magnitude lower than erg s-1.
@@
The Large Magellanic Cloud (LMC) represents an ideal laboratory to hunt for stellar mass BHs with the above techniques owing to its homogeneous sample of well-characterised massive stars and low interstellar extinction compared to the Milky Way (Evans et al. 2011; Sana et al. 2013; Evans et al. 2015; Villaseñor et al. 2021; Mahy et al. 2020b; Shenar et al. 2022b). Several single-lined spectroscopic (SB1) binaries have been identified (Villaseñor et al. 2021; Shenar et al. 2022b). The Chandra T-ReX programme (Crowther et al. 2022) has shown that faint X-ray sources down to 1031 erg s-1 can be detected, but the statistics of X-ray-faint (1031-1035 erg s-1) BHs is still unclear. We do a population synthesis study to predict the X-ray luminosity function of BH+OB binaries in the LMC with specific attention to the faint end of the distribution. We discuss the possibility of confirming the presence of a candidate BH through targeted X-ray searches with Chandra (Evans & Civano 2018), and SRG/eROSITA (Predehl et al. 2021; Sunyaev et al. 2021).
@@
Section 2 investigates the radiative mechanisms around BHs accreting at low mass accretion rates. Section 3 describes the detailed binary evolution models and our assumptions to estimate the X-ray luminosity during the BH+OB star phase from the models. We describe our results in Sect. 4 and compare our model predictions with observations in Sect. 5. We discuss relevant uncertainties and observational challenges in Sect. 6. Section 7 summarises the takeaway messages from our work and future directions.
@@ 2 Radiative efficiency
When a black hole accretes plasma at a rate (in yr-1, measured within a few times the event horizon’s radius), the release of gravitational potential energy is susceptible to produce X-rays with a luminosity , with the radiative efficiency and is the speed of light in vacuum. The radiative efficiency of this plasma (and, in turn, the spectral properties of the emission) depends on the flow geometry and the mass accretion rate (Liu & Qiao 2022). If the plasma is supplied by a Roche lobe overflowing stellar companion, a disk unavoidably forms (Savonije 1978). If the companion to the BH is not Roche-lobe filling, the flow geometry is set by the specific angular momentum of the accreted plasma from the stellar wind of the companion (Sect. 3.2).
@@
In massive systems where the stellar companion to the BH fills its Roche lobe, the high mass transfer rate may lead to a super-Eddington accretion ( ¿ 1) and the formation of ultraluminous X-ray binaries (e.g., Quast et al. 2019). When the stellar companion to the BH does not fill its Roche lobe, the BH can accrete mass from the stellar wind of the companion. In this case, an accretion disk can form if the specific angular momentum of the accreted stellar wind matter is sufficient to orbit around the BH beyond its event horizon.
@@ 2.1 Sub-Eddington accretion regime
When an accretion disk forms around a BH with a stellar companion that does not fill its Roche lobe, the mass accretion rate typically lies between 0.1% and 100% of the Eddington mass accretion rate (e.g., see Fig. A.1 of Sen et al. 2021). Accretion is mediated by a geometrically thin and optically thick accretion disk. The radiative efficiency lies between and , depending on the black hole spin and the misalignment between the disk and the black hole spin (Novikov & Thorne 1973).
@@
A decrease in the mass accretion rate without a change in the accretion flow geometry is possible, for instance, if the density of the OB star wind drops without changes in the wind velocity. In this case, the plasma might no longer be dense enough to cool radiatively at a rate which balances viscous heating. The disk thickens and becomes a ‘slim-disk’, as the relative contribution of the thermal support increases. A large fraction of the gravitational potential energy released might be advected into the black hole’s event horizon, as described by the ADAF (Ichimaru 1977; Narayan & Yi 1994). Therefore, the radiative efficiency drops. It might be the reason for the transition between the high/soft and the low/hard states we observe in black hole hosting X-ray binaries (Belloni 2010). It could also explain the low X-ray luminosity of the BH candidate around the O-type main-sequence star HD96670, in spite of its short orbital period (Gomez & Grindlay 2021).
@@
Alternatively, if the flow is magnetized, it can also transit to a magnetically-arrested disk (MAD, Begelman et al. 2022) where accretion proceeds through episodic ejection of magnetic bubbles via the interchange instability (Porth et al. 2021) and through a reconnecting current sheet between the inner edge of the disk and the black hole event horizon (Ripperda et al. 2022). On the other hand, if the mass supply rate decreases due to an increase in wind speed and/or orbital separation, it is accompanied by a drop in specific angular momentum and the disk vanishes.
@@
Owing to the scale-invariant properties of black holes, emission from supermassive black holes can provide fruitful insights to model X-ray faint stellar-mass black holes. Sagittarius A∗ is thought to be fed by the winds from orbiting nearby Wolf-Rayet stars (Russell et al. 2017; Ressler et al. 2023). Only a fraction of the mass supplied at the Bondi radius reaches the black hole event horizon though: Faraday rotation (Bower et al. 2003; Marrone et al. 2007; Wang et al. 2013) and extrapolation of magneto-hydrodynamic simulations (Ressler et al. 2018) lead to in the immediate vicinity of the BH, but the presence of a disk is unclear. In spite of its status of active galactic nucleus, M87∗ also accretes at a low rate, with (Prieto et al. 2016), orders of magnitude below the geometrically-thin, optically-thick disk regime.
@@
It was found from polarimetry that the emission from these objects originates in a dilute and highly magnetized corona populated with hot synchrotron-emitting electrons (Bower et al. 2018). Similar conclusions were drawn for stellar-mass black holes accreting at a rate below (e.g. Cygnus X-1 in the low/hard state, Cangemi et al. 2021). The range of mass accretion rates onto stellar mass BHs from the stellar wind of its companion is also expected to be the range - 10-2 (Fig. A.1 of Sen et al. 2021). In both stellar-mass and supermassive BHs, coronal heating is ensured by particle acceleration processes at play once the plasma becomes dilute enough to be collisionless (Gruzinov 1998).

### v2

@@ Abstract
Context. Recently, astrometric and spectroscopic surveys of OB stars revealed a few stellar-mass black holes (BHs) with orbital periods of as low as 10 days. Contrary to wind-fed BH high-mass X-ray binaries (HMXBs), no X-ray counterpart was detected, probably because of the absence of a radiatively efficient accretion disc around the BH. Nevertheless, dissipative processes in the hot, dilute, and strongly magnetised plasma around the BH (so-called BH corona) can still lead to non-thermal X-ray emission (e.g. synchrotron).
@@
Methods. We used detailed binary evolution models computed with MESA for initial primary masses of 10-90 and orbital periods of 1-3000 d. We computed the X-ray luminosity for a broad range of radiative efficiencies that depend on the mass accretion rate and flow geometry.
Results. For typical conditions around stellar-mass BHs, we show that particle acceleration through magnetic reconnection can heat the BH corona. A substantial fraction of the gravitational potential energy from the accreted plasma is converted into non-thermal X-ray emission. Our population synthesis analysis predicts that at least 28 (up to 72) BH+OB star binaries in the Large Magellanic Cloud (LMC) produce X-ray luminosities of above 1031 erg s-1, which are observable with focused Chandra observations. We identify a population of SB1 systems in the LMC and HD96670 in the Milky Way comprising O stars with unseen companions of masses of above 2.3 , which aligns well with our predictions and may be interesting sources for follow-up observations. The predicted luminosities of the OB companions to these X-ray-emitting BHs are 104.5-5.5 .
Conclusions. These findings advocate for prolonged X-ray observations of the stellar-mass black hole candidates identified in the vicinity of OB stars. Such long exposures could reveal the underlying population of X-ray-faint BHs and provide constraints for the evolution from single to double degenerate binaries and identify the progenitors of gravitational wave mergers.
@@
The detection of merging stellar-mass black holes (BHs) and neutron stars (Abbott et al. 2016, 2019) has triggered a thrilling quest to discover their progenitors. Co-evolution of massive binary stars (i.e. hydrogen burning starts at the same instant of time in both binary components) is a possible formation channel of merging compact objects (Belczynski et al. 2008; Marchant et al. 2016; Kruckow et al. 2018; Spera et al. 2019; Vigna-Gómez et al. 2019; Mapelli 2020; Bavera et al. 2021; Marchant et al. 2021; Jiang et al. 2023). The stable mass-transfer channel has been shown to significantly contribute to double compact object mergers (van den Heuvel et al. 2017; Gallegos-Garcia et al. 2021; van Son et al. 2022; Shao & Li 2022; Briel et al. 2023; Picco et al. 2024; Dorozsmai & Toonen 2024; Olejak et al. 2024). However, many unconstrained assumptions (such as mass and angular momentum loss, internal mixing, and stellar winds) used in the modelling of binary evolution cripple our capacity to make reliable predictions regarding the rates and properties of double compact object mergers (Broekgaarden et al. 2022; Mandel & Broekgaarden 2022; Marchant & Bodensteiner 2023).
@@
In the co-evolutionary scenario, high-mass X-ray binaries (HMXBs) represent an ephemeral albeit decisive phase before the formation of double compact objects from massive binaries. HMXBs contain a massive star in orbit with an X-ray-bright compact object (Walter et al. 2015; Kretschmar et al. 2019; Motta et al. 2021; Chaty 2022; Fortin et al. 2023). A few of them were found to host stellar-mass BHs accreting stellar material from mass transfer via Roche lobe overflow and/or the capture of a fraction of the stellar wind (e.g. Cygnus X-1 Orosz et al. 2011, and LMC X-1 Orosz et al. 2009). Beyond their role in explaining mergers, these systems are excellent test beds for theories of stellar evolution (Belczynski et al. 2008; Fishbach & Kalogera 2022), gravity and magnetism in the strong field regime (Narayan & Yi 1995; Bozzo et al. 2008; Karino et al. 2019), stellar winds, and X-ray irradiation (Krtička et al. 2018; Vilhu et al. 2021).
@@
The plasma surrounding accreting stellar-mass BHs is a source of X-rays whose intensity and spectral properties depend on the flow geometry (disc or sphere-like) and on the mass-accretion rate (Shakura 1973; Narayan & Yi 1995; Frank et al. 2002). At high mass-accretion rates, a geometrically thin and optically thick accretion disc can form that is radiatively efficient and X-ray bright. However, at lower mass-accretion rates (Sharma et al. 2007; Xie & Yuan 2012) and/or when the flow is not centrifugally supported (Narayan & Yi 1994), an accretion disc seldom forms and the radiative efficiency of the plasma drops. In very sub-Eddington accretion flows, most of the kinetic energy of the electrons is not radiated away but is advected into the BH event horizon (ADAF; Ichimaru 1977; Narayan & Yi 1994; Narayan et al. 1998; Yuan 2001). Also, a significant fraction of the material captured by the gravitational potential of the BH may not eventually accrete onto the BH, instead being lost in strong winds (ADIOS – Blandford & Begelman 1999; Yuan et al. 2012, 2015). As a result, the X-ray luminosity of most stellar-mass BHs might be below the detectability threshold of all-sky instruments ( erg s-1, taken from Vanbeveren et al. 2020), while remaining detectable during outbursts.
@@
Population synthesis models predict thousands of BHs in orbit with a main sequence, giant, or supergiant companion in the Milky Way (Shao & Li 2019, 2020). Nevertheless, most BHs in a binary with a main sequence OB star are expected to be X-ray faint due to the inability to form a radiatively efficient accretion disc around the BH (Sen et al. 2021; Hirai & Mandel 2021). In the absence of an accretion disc, X-rays are dominated by non-thermal emission from electrons accelerated in the hot, dilute, magnetised and spherical (i.e. non-centrifugally maintained) region around the BH, called the BH ‘corona’ (Bisnovatyi-Kogan & Lovelace 1997; Quataert & Narayan 1999). This radiative model has been corroborated by observations of supermassive BHs like Sagittarius A∗ (Baganoff et al. 2003), M87∗ (Event Horizon Telescope Collaboration et al. 2019), and NGC 3115 (Wong et al. 2011). In this work, we use the above model to investigate the properties of X-ray emission from stellar-mass BHs in orbit with a main sequence OB star.
@@
The sample of detected stellar-mass BHs is likely very incomplete. The Milky Way and Magellanic Clouds are predicted to harbour millions and hundreds of stellar-mass BHs, respectively (Timmes et al. 1996; Breivik et al. 2017; Langer et al. 2020; Chawla et al. 2022; Jayasinghe et al. 2023). Nevertheless, only a few tens of these objects have been confirmed through X-ray detection (Walter et al. 2015; Corral-Santana et al. 2016), with a further handful recently confirmed through monitoring of the stellar companion (Giesers et al. 2018; Thompson et al. 2019; Liu et al. 2019; Rivinius et al. 2020; Masuda & Hirano 2021; Jayasinghe et al. 2021; Gomez & Grindlay 2021; Saracino et al. 2022; Shenar et al. 2022a; Tanikawa et al. 2023; El-Badry et al. 2023a, b; Gaia Collaboration et al. 2024). Most of the approximately 60 stellar-mass BHs detected through X-ray emission are transients, with periodicities of the order of years to decades (McClintock & Remillard 2006; Corral-Santana et al. 2016). In contrast, their X-ray luminosity during quiescence is several orders of magnitude lower than erg s-1.
@@
The Large Magellanic Cloud (LMC) represents an ideal laboratory in which to hunt for stellar mass BHs with the above techniques owing to its homogeneous sample of well-characterised massive stars and low interstellar extinction compared to the Milky Way (Evans et al. 2011; Sana et al. 2013; Evans et al. 2015; Villaseñor et al. 2021; Mahy et al. 2020b; Shenar et al. 2022b). Several single-lined spectroscopic (SB1) binaries have been identified (Villaseñor et al. 2021; Shenar et al. 2022b). The Chandra T-ReX programme (Crowther et al. 2022) has shown that faint X-ray sources down to 1031 erg s-1 can be detected, but the statistics regarding X-ray-faint (1031-1035 erg s-1) BHs are still unclear. Here, we present a population synthesis study to predict the X-ray luminosity function of BH+OB binaries in the LMC with specific attention to the faint end of the distribution. We discuss the possibility of confirming the presence of a candidate BH through targeted X-ray searches with Chandra (Evans & Civano 2018) and SRG/eROSITA (Predehl et al. 2021; Sunyaev et al. 2021).
@@
In Section 2, we present an investigation of the radiative mechanisms around BHs accreting at low mass-accretion rates. In Section 3, we describe the detailed binary evolution models and the assumptions we make when estimating the X-ray luminosity during the BH+OB star phase from the models. We describe our results in Sect. 4 and compare our model predictions with observations in Sect. 5. We discuss relevant uncertainties and observational challenges in Sect. 6. Section 7 summarises the takeaway messages from our work and future directions.
@@ 2 Radiative efficiency
When a black hole accretes plasma at a rate (in yr-1, measured within a few times the event horizon’s radius), the release of gravitational potential energy is susceptible to produce X-rays with a luminosity , with the radiative efficiency and is the speed of light in vacuum. The radiative efficiency of this plasma (and, in turn, the spectral properties of the emission) depends on the flow geometry and the mass accretion rate (Liu & Qiao 2022). If the plasma is supplied by a Roche lobe overflowing stellar companion, a disc unavoidably forms (Savonije 1978). If the companion to the BH is not Roche-lobe filling, the flow geometry is set by the specific angular momentum of the accreted plasma from the stellar wind of the companion (Sect. 3.2).
@@
In massive systems where the stellar companion to the BH fills its Roche lobe, the high mass transfer rate may lead to a super-Eddington accretion ( ¿ 1) and the formation of ultraluminous X-ray binaries (e.g. Quast et al. 2019). When the stellar companion to the BH does not fill its Roche lobe, the BH can accrete mass from the stellar wind of the companion. In this case, an accretion disc can form if the specific angular momentum of the accreted stellar wind matter is sufficient to orbit around the BH beyond its event horizon.
@@ 2.1 Sub-Eddington accretion regime
When an accretion disc forms around a BH with a stellar companion that does not fill its Roche lobe, the mass accretion rate typically lies between 0.1% and 100% of the Eddington mass accretion rate (e.g. see Fig. A.1 of Sen et al. 2021). Accretion is mediated by a geometrically thin and optically thick accretion disc. The radiative efficiency lies between and , depending on the black hole spin and the misalignment between the disc and the black hole spin (Novikov & Thorne 1973).
@@
A decrease in the mass accretion rate without a change in the accretion flow geometry is possible, for instance, if the density of the OB star wind drops without changes in the wind velocity. In this case, the plasma might no longer be dense enough to cool radiatively at a rate which balances viscous heating. The disc thickens and becomes a ‘slim-disc’, as the relative contribution of the thermal support increases. A large fraction of the gravitational potential energy released might be advected into the black hole’s event horizon, as described by the ADAF (Ichimaru 1977; Narayan & Yi 1994). Therefore, the radiative efficiency drops. It might be the reason for the transition between the high/soft and the low/hard states we observe in black hole hosting X-ray binaries (Belloni 2010). It could also explain the low X-ray luminosity of the BH candidate around the O-type main-sequence star HD96670, in spite of its short orbital period (Gomez & Grindlay 2021).
@@
Alternatively, if the flow is magnetised, it can also transit to a magnetically arrested disc (MAD, Begelman et al. 2022) where accretion proceeds through episodic ejection of magnetic bubbles via the interchange instability (Porth et al. 2021) and through a reconnecting current sheet between the inner edge of the disc and the black hole event horizon (Ripperda et al. 2022). On the other hand, if the mass supply rate decreases due to an increase in wind speed and/or orbital separation, it is accompanied by a drop in specific angular momentum and the disc vanishes.
@@
Owing to the scale-invariant properties of black holes, emission from supermassive black holes can provide fruitful insights to model X-ray faint stellar-mass black holes. Sagittarius A∗ is thought to be fed by the winds from orbiting nearby Wolf-Rayet stars (Russell et al. 2017; Ressler et al. 2023). Only a fraction of the mass supplied at the Bondi radius reaches the black hole event horizon though: Faraday rotation (Bower et al. 2003; Marrone et al. 2007; Wang et al. 2013) and extrapolation of magneto-hydrodynamic simulations (Ressler et al. 2018) lead to in the immediate vicinity of the BH, but the presence of a disc is unclear. In spite of its status of active galactic nucleus, M87∗ also accretes at a low rate, with (Prieto et al. 2016), orders of magnitude below the geometrically thin, optically thick disc regime.
@@
It was found from polarimetry that the emission from these objects originates in a dilute and highly magnetised corona populated with hot synchrotron-emitting electrons (Bower et al. 2018). Similar conclusions were drawn for stellar-mass black holes accreting at a rate below (e.g. Cygnus X-1 in the low/hard state, Cangemi et al. 2021). The range of mass accretion rates onto stellar mass BHs from the stellar wind of its companion is also expected to be the range - 10-2 (Fig. A.1 of Sen et al. 2021). In both stellar-mass and supermassive BHs, coronal heating is ensured by particle acceleration processes at play once the plasma becomes dilute enough to be collisionless (Gruzinov 1998).